OpenStudy (anonymous):

Compare the function with the parent function. Without graphing, what are the vertex, axis of symmetry, and transformations of the parent function? A)(–1/4, –7); x = –1/4; translated to the right 1/4 units and up 7 units B)(1/4, 7); x = 1/4; translated to the right 1/4 units and up 7 units C)(–1/4, 7); x = –1/4; translated to the left 1/4 units and up 7 units D)(1/4, –7); x = 1/4; translated to the right 1/4 units and down 7 units

OpenStudy (anonymous):

@tyteen4a03 i think its A but im not 100% sure.

OpenStudy (tyteen4a03):

This looks like 4 separate questions... doesn't look like MC answers to me.

OpenStudy (anonymous):

this is the question: Compare the function with the parent function. Without graphing, what are the vertex, axis of symmetry, and transformations of the parent function? these are the choices i can choose from: A)(–1/4, –7); x = –1/4; translated to the right 1/4 units and up 7 units B)(1/4, 7); x = 1/4; translated to the right 1/4 units and up 7 units C)(–1/4, 7); x = –1/4; translated to the left 1/4 units and up 7 units D)(1/4, –7); x = 1/4; translated to the right 1/4 units and down 7 units

OpenStudy (tyteen4a03):

Where is the parent function?

OpenStudy (anonymous):

y = │8x – 2│ – 7 sorry

OpenStudy (anonymous):

@tyteen4a03

OpenStudy (tyteen4a03):

Still looking... hang on.

OpenStudy (tyteen4a03):

The transformations mentioned are shifting graphs left/right and up/down. Remember that f(x) + C moves the graph up (negative number moves it down) and f(x + C) moves it left (negative moves it right). Try reforming the equations according to these rules then see where the vertex lies.

OpenStudy (anonymous):

the lines on the right. am i right with A as my answer?